OpenStudy (selena2345):

If 3 pounds of apples cost $0.90,how much will 10 pounds cost ?

OpenStudy (anonymous):

10 lbs would cost $3.00.

OpenStudy (darkbluechocobo):

^ yus

OpenStudy (anonymous):

Sorry I shouldn't of had given you the answer. I can tell you how to do it if you'd like.

OpenStudy (selena2345):

ok @nickmifsud1223

OpenStudy (tkhunny):

1) Calculate a unit rate. $0.90/(3 apple) = 30¢ / apple. Now, how much for 10 apples? 2) Set up a proportion. \(\dfrac{90¢}{3\;apples} = \dfrac{What}{10\;apples}\) 3) Can you think of another way?

OpenStudy (anonymous):

That's how you do it. ^

OpenStudy (texaschic101):

I agree....I myself would probably do it the proportion way

OpenStudy (selena2345):



OpenStudy (anonymous):

Yeah the proportion way is easier to me. But whatever works best.

OpenStudy (tkhunny):

Just remember you are working in pennies. 3) Set up a function! After calculating the unit rate, Price(Apples) = 0.30*Apples. Substitute the desired value and you are done.